# Learning Diffusion Policies for Robotic Manipulation of Timber Joinery under Fabrication Uncertainty

### v1

@@ Abstract
Construction uncertainties such as fabrication inaccuracies and material imperfections pose a significant challenge to contact-rich robotic manipulation by hindering precise and robust assembly. In this paper, we explore the performance and robustness of diffusion policy learning as a promising solution for contact-sensitive robotic assembly at construction scale, using timber mortise and tenon joints as a case study. A two-phase study is conducted: first, to evaluate policy performance and applicability; second, to assess robustness in handling fabrication uncertainties simulated as randomized perturbations to the mortise position. The best-performing policy achieved a total average success rate of 75% with perturbations up to 10 mm, including 100% success in unperturbed cases. The results demonstrate the potential of sensory–motor diffusion policies to generalize to a wide range of complex, contact-rich assembly tasks across construction and manufacturing, advancing robotic construction under uncertainty and contributing to safer, more efficient building practices.
@@
Contact-rich manipulation, Robotic assembly, Fabrication uncertainty, Diffusion policy, Timber joinery, Robot learning, Construction robotics
@@ 1 Introduction
Construction plays a critical role in the global economy but continues to face long-standing challenges, including stagnant productivity growth, shortages of skilled labor, and persistent health and safety concerns Delgado et al. [2019], Wei et al. [2023]. The industry often requires workers to perform repetitive and physically demanding tasks, such as lifting and positioning heavy components, which can contribute to chronic injuries and reduced long-term workforce capacity Laukkanen [1999], Arndt et al. [2005]. Automation, and in particular robotic assembly, offers the potential to address these issues by enhancing precision, increasing efficiency, enabling mass customization, and reducing the physical burden on workers Bock [2015], Musarat et al. [2024], Chen and Adel [2025]. Recent advances in construction robotics have increasingly harnessed the precision and programmability of industrial robotic arms to execute complex and customized assemblies that would be difficult to achieve through manual methods Graser et al. [2021], Chen and Adel [2025]. Despite these advancements, translating the capabilities of industrial robots from manufacturing to the uncertain and unstructured conditions of construction, characterized by fabrication inaccuracies, material imperfections, as well as dynamic and evolving environments, remains a significant challenge, making it difficult to achieve the same levels of repeatability, quality, and precision that are standard in factory settings Adel et al. [2024], Delgado et al. [2019], Musarat et al. [2024].
@@
The momentum in construction robotics research is especially evident in recent work on robotic timber assembly Willmann et al. [2017], Apolinarska [2018], Adel et al. [2018], Adel [2020, 2022], Chai et al. [2022], Lauer et al. [2023], where many workflows rely on planar face-to-face or butt joints, whose geometric simplicity allows the use of pre-programmed trajectories for successful assembly. In contrast, the manipulation of traditional timber joinery such as mortise and tenon, lap, or scarf joints requires contact-sensitive insertion and interlocking, demanding high precision and control of contact forces Apolinarska [2018], Leung et al. [2021]. Similar to many contact-rich robotic manipulation tasks Suomalainen et al. [2022], the assembly of these joints can lead to misalignment and jamming if not coupled with real-time control for robust error recovery Apolinarska et al. [2021], Kramberger et al. [2022]. As a result, these contact-sensitive tasks are often delegated to skilled human workers, who intervene to clamp, hammer, and adjust fits, compensating for fabrication inaccuracies and material imperfections Yang et al. [2024]. While effective in the short term, such reliance on manual intervention can break the continuity of digital production workflows, introducing delays and reducing overall efficiency. Furthermore, it increases reliance on a shrinking pool of skilled labor and entails physically demanding, labor-intensive work that can contribute to worker fatigue and injury.
@@
Figure 1: Overview of our method. A CNN-based diffusion policy is trained conditional on end-effector pose and F/T data observations to predict sequences of robot actions.
@@
Timber joinery is a centuries-old construction technique of geometrically interlocking timber components to create durable structural connections with minimal reliance on fasteners or adhesives Benson [1981]. These joints are valued for their enhanced structural performance, rotational stiffness, and resilience under seismic and fire conditions Fang and Mueller [2023], Fang [2021]. While they offer significant structural advantages, their complex geometry and contact-rich assembly make them particularly sensitive to the uncertainties inherent to construction tasks, including fabrication inaccuracies from cut length and angle deviations and component shifting during picking, placing, and fastening, as well as material imperfections from dimensional variations and hygrothermal expansion Adel et al. [2024]. Collectively, these factors further increase the complexity of automating the assembly of such joints in construction settings. Despite the significance of this challenge, automated assembly of traditional joinery remains nascent in robotic workflows within the Architecture, Engineering, and Construction (AEC) domain, particularly when accounting for fabrication and material uncertainty. Overcoming these limitations requires feedback-driven and closed-loop planning and control methods capable of adapting to unpredictable conditions inherent in construction workflows.
@@
Learning-based methods have increasingly replaced traditional control-based methods for planning robot actions. Sensory–motor policy learning methods integrate sensor observations, planning, and control into a single end-to-end system, originally implemented using deep convolutional neural networks (CNNs) Levine et al. [2016]. These methods enable the robot to learn how to act in response to what it senses without requiring separately engineered modules for perception, planning, state estimation, or low-level control. These learned control policies map raw observation inputs, such as force feedback and images captured from a camera, directly to motor commands or robot actions using various learning-based methods Finn et al. [2017], Levine et al. [2018], Kalashnikov et al. [2018], Brohan et al. [2023]. Among these methods, diffusion policy learning via behavior cloning Chi et al. [2024a] has shown outstanding performance for various manipulation scenarios Chi et al. [2024b], Hou et al. [2025], Yang et al. [2025], Zhu et al. [2025]. The key advantages of diffusion policies lie in their ability to learn multi-modal action distributions, capturing multiple plausible ways to perform a task, while maintaining training stability and predicting coherent sequences of future actions Chi et al. [2024a].
@@
The research presented in this paper tackles the challenge of contact-rich robotic manipulation under construction uncertainties, where fabrication inaccuracies and material imperfections hinder precise and robust assembly. We address this by implementing diffusion policy learning, trained on robot pose and force/torque (F/T) data from teleoperated demonstrations, to plan and control the construction-scale assembly of mortise and tenon joints as a case study (Fig. 1).
1.1 Related work
@@
In construction with discrete elements, such as brickwork and timber or steel framing, uncertainty remains a fundamental challenge to achieving precise and robust robotic assembly. In timber framing, which serves as the case study in this research, these uncertainties arise from material imperfections (e.g., dimensional deviations, deformation, shrinkage/expansion due to moisture) and fabrication inaccuracies (e.g., cut length and angle errors, pose deviations from calibration or handling). If unaccounted for, these variabilities accumulate during assembly, leading to misalignments, part collisions, and robot failures when trajectories are executed directly from as-planned digital models  Adel et al. [2024]. Prior research, in robotic timber assembly specifically, has explored various strategies to address these challenges, including the development of an external tracking system for end-effector pose correction in multi-robot assembly workflows Stadelmann et al. [2019], computational modeling of tolerance propagation Gandia et al. [2022], and adaptive, feedback-driven methods Helm et al. [2016], Adel et al. [2024], Ruan et al. [2023], Cote et al. [2024]. The successful and precise execution of timber joints often still depends on human intervention to correct deviations during nailing, gluing, and insertions Willmann et al. [2017], Adel et al. [2018], Adel [2020], Helmreich et al. [2022], Adel et al. [2024]. The challenges of uncertainty become even more pronounced in contact-sensitive assembly tasks, such as those for traditional timber joinery, where success depends not only on geometric precision but also on the ability to control excessive force interactions imposed by friction and tight tolerances. In such scenarios, even minor deviations introduced during fabrication or caused by material imperfections can easily lead to misalignment, resulting in jamming, stick-slip behavior, and material damage Apolinarska et al. [2021].
@@
Handling force interactions in contact-rich robotic manipulation is often addressed through model-based impedance or admittance control laws Albu-Schäffer et al. [2007], Suárez-Ruiz et al. [2018], learning-based control policies Vecerik et al. [2019], Schoettler et al. [2020] or hybrid approaches combining both Johannsmeier et al. [2019], Hou et al. [2025]. In the context of assembly with timber joinery, a few studies have explored learning-based methods; Apolinarska et al. applied reinforcement learning for the motion control of an industrial robot to assemble single- and double-lap joints using robot pose and force/torque data from a wrist-mounted sensor Apolinarska et al. [2021]. For the single-lap tasks, they also investigated the policy robustness by introducing angular and linear offsets to the initial robot pose, grasping the top piece. Their results showed strong performance in zero-offset cases and revealed challenges of policy adaptation when angular and linear offsets, as well as angled or double insertions, were introduced. The findings underscored the importance of robust policy learning for handling geometric variations and positional deviations in contact-rich assembly tasks. Kramberger et al. also proposed a human–robot collaborative workflow using a learning-by-demonstration method that integrated a compliance controller into Cartesian-space dynamic movement primitives for perpendicular single-lap joint insertion; however, they relied on significantly loose tolerances Kramberger et al. [2022].
@@
Other studies have explored semi-structured or open-loop methods for timber joinery assembly; Robeller et al. introduced a custom end-effector that generated vibration combined with a manually applied pulse force using a mallet to facilitate the connection of timber panels using dovetail joints Robeller et al. [2017]. Leung et al. demonstrated the use of an industrial robot in combination with manually placed distributed robotic clamps to assemble tight-fitting half-lap joints, highlighting the potential for precise and repeatable assembly. During their experiments, they observed challenges such as misalignments and collisions—largely arising from fabrication inaccuracies and deviations between digital and as-built models—that are common in construction-scale robotic applications Leung et al. [2021]. Rogeau employed visual feedback and fiducial markers for the assembly of wood panels with multiple mortise and tenon joints and reported successful full insertions in about 50% of their test cases across various joint configurations such as mortise tightness and tenon chamfer angles Rogeau [2023].
@@
Diffusion policies Chi et al. [2024a], Ze et al. [2024], Zhao et al. [2025] represent a recent class of behavior cloning methods built upon Denoising Diffusion Probabilistic Models (DDPMs) Ho et al. [2020], Sohl-Dickstein et al. [2015] and Denoising Diffusion Implicit Models (DDIMs) Song et al. [2022]. DDPMs are generative models originally developed for high-quality image synthesis, transforming random noise into samples from a target distribution through an iterative denoising process. DDIMs extend this framework by formulating a non-Markovian reverse process, which allows the number of denoising steps used during inference to be reduced without changing the training procedure, thereby enabling significantly faster sampling. In the diffusion policy via robot action diffusion Chi et al. [2024a], the denoising process generates a sequence of robot trajectories at each inference step. The approach is known to naturally capture multimodal action distributions, meaning it can represent multiple distinct yet equally valid action sequences for completing a task, as often occurs in human demonstrations. The approach also maintains temporal coherence, meaning the policy predicts an entire sequence of future actions jointly rather than one step at a time. This joint prediction ensures that consecutive actions remain consistent (e.g., following one continuous path) rather than switching erratically between different predicted single actions, which can lead to unstable or jittery motion. Moreover, the method avoids the need for action discretization (breaking continuous movements into a fixed set of bins) or negative sampling (generating incorrect actions during training to guide learning) Chi et al. [2024a]. This removes common limitations of explicit behavior cloning methods Florence et al. [2020], Shafiullah et al. [2022], which can make it difficult to achieve high-precision actions. Diffusion policies also offer improved training stability compared to implicit behavior cloning methods Florence et al. [2022], Jarrett et al. [2020]. In addition, they have demonstrated robustness to perturbations, visual occlusions, and idle actions in some manipulation tasks Chi et al. [2024a].
Planning and control through contact remains a major challenge for both model- and learning-based approaches, due to the inherent complexity of contact dynamics Yang et al. [2025]. While diffusion policies have achieved notable success in various dexterous manipulation tasks, their effectiveness in contact-rich scenarios is still under active investigation Hou et al. [2025], Kang et al. [2025], Wu et al. [2025]. Additionally, despite significant progress in construction robotic research, the application of robot learning for planning and control, particularly for contact-sensitive tasks, remains underexplored. In this study, we evaluate the performance and robustness of diffusion policies through a case study on the assembly of timber mortise and tenon joints under fabrication uncertainty. The following section outlines the objectives, novelty, and contributions of this research.
1.2 Objectives and contributions
In the research presented in this paper, we develop a sensory–motor diffusion policy learning framework via behavior cloning for the contact-rich robotic manipulation at construction scale. The study pursues three primary objectives: (1) Demonstrate applicability and performance of diffusion policies through a case study on the robotic assembly of mortise and tenon timber joints; (2) Investigate policy robustness under construction-relevant uncertainties, simulated as randomized perturbations to the mortise position; and (3) Identify actionable guidelines for data collection strategies and policy parameter tuning that yield the most successful and robust performance, providing a practical reference for future research and deployment.
This paper presents one of the first systematic investigations in the AEC domain applying diffusion policy learning to robot planning and control. By enabling autonomous timber joinery assembly without reliance on hammering or human intervention, the proposed framework addresses physically demanding and craftsmanship-intensive tasks, offering a scalable learning framework for contact-sensitive manipulation tasks. The approach demonstrates potential for generalizing a wide spectrum of contact-rich assembly challenges, including complex timber joints, pipe fitting, and light metal framing. Ultimately, this work contributes to advancing robotic construction under uncertainty, bridging state-of-the-art generative policy learning with real-world AEC workflows, and provides both theoretical and practical advances that can accelerate adoption across the field.

### v2

@@ Abstract
Fabrication uncertainties, such as tolerance accumulation and material imperfections, pose a significant challenge to contact-rich robotic manipulation in construction by hindering precise and robust assembly. In this paper, we investigate the performance and robustness of diffusion policy learning for contact-rich assembly at the construction scale, using a tight-fitting timber mortise and tenon joint as a case study. A two-phase experimental study is conducted: first, to evaluate baseline policy performance and applicability; second, to assess policy robustness under fabrication-induced uncertainties modeled as randomized perturbations to the mortise position. The diffusion policy is trained on teleoperated demonstrations using an industrial robotic arm conditioned on end-effector pose and force/torque feedback. The best-performing policy achieved a total average success rate of 75% under perturbations up to 10 mm, including 100% success in unperturbed cases. The results demonstrate the potential of sensory–motor diffusion policies to enable high-precision contact-rich manipulation on large-scale industrial robotic arms, reducing reliance on skilled manual intervention. This work advances robotic construction under uncertainty and provides practical insights for deploying learning-based control in real-world Architectural, Engineering, and Construction applications.
@@
Contact-rich manipulation, Diffusion policy, Fabrication uncertainty, Timber joinery, Robotic assembly, Construction robotics
@@ 1 Introduction
Construction plays a critical role in the global economy but continues to face long-standing challenges, including stagnant productivity growth, shortages of skilled labor, and persistent health and safety concerns Delgado et al. [2019], Wei et al. [2023]. The industry often requires workers to perform repetitive, physically demanding tasks, such as lifting and positioning heavy components, which can lead to chronic injuries and reduced long-term workforce capacity Laukkanen [1999], Arndt et al. [2005]. Automation, and in particular robotic assembly, offers the potential to address these issues by enhancing precision, increasing efficiency, enabling mass customization, and reducing the physical burden on workers Bock [2015], Musarat et al. [2024]. Recent advances in construction robotics have increasingly utilized industrial robotic arms, which play a central role in large-scale robotic assembly due to their high payload capacity and long reach to execute customized and complex assemblies that would be difficult to achieve manually Chen and Adel [2025]. Despite these advancements, translating the capabilities of industrial robots from manufacturing to the uncertain and unstructured conditions of construction, characterized by fabrication inaccuracies, material imperfections, as well as dynamic and evolving environments, remains a significant challenge, making it difficult to achieve the same levels of repeatability, quality, and precision that are standard in factory settings Adel et al. [2024], Delgado et al. [2019], Musarat et al. [2024].
@@
The momentum in construction robotics research is especially evident in recent work on robotic timber assembly Willmann et al. [2017], Apolinarska [2018], Adel et al. [2018], Adel [2020], Graser et al. [2021], Chai et al. [2022], Lauer et al. [2023], where many workflows rely on planar face-to-face or butt joints, whose geometric simplicity allows the use of preprogrammed trajectories for successful assembly. In contrast, the manipulation of timber joinery such as mortise and tenon, lap, or scarf joints requires contact-rich insertion and interlocking, demanding high precision and control of contact forces Apolinarska [2018], Leung et al. [2021]. In these scenarios, misalignment and jamming can occur if insertion is not coupled with real-time control and robust error recovery Apolinarska et al. [2021], Kramberger et al. [2022]. As a result, these tasks are often delegated to skilled human workers, who intervene to clamp, hammer, and adjust fits, compensating for fabrication inaccuracies and material imperfections Yang et al. [2024]. While effective in the short term, such reliance on manual intervention can break the continuity of digital production workflows, introducing delays and reducing overall efficiency. Furthermore, it increases reliance on a shrinking pool of skilled labor and entails physically demanding, labor-intensive work that can contribute to worker fatigue and injury.
@@
Figure 1: Overview of our method. A CNN-based diffusion policy is trained conditional on end-effector pose (i.e., position and orientation) and F/T observations to predict sequences of robot actions.
@@
Timber joinery is a centuries-old construction technique of geometrically interlocking timber components to create durable structural connections with minimal reliance on fasteners or adhesives Benson [1981]. These joints are valued for their enhanced structural performance, rotational stiffness, and resilience under seismic and fire conditions Fang and Mueller [2023], Fang [2021]. However, their complex geometry and contact-dominated assembly process make them particularly sensitive to fabrication inaccuracies and material imperfections, including cut length and angle deviations, component shifting during picking, placing, and fastening, and dimensional changes due to hygrothermal effects Adel et al. [2024]. These factors significantly complicate robotic automation of joinery assembly. Despite the significance of this challenge, automated assembly of traditional joinery remains nascent in robotic workflows within the Architecture, Engineering, and Construction (AEC) domain, particularly when explicitly accounting for fabrication-induced and material-driven variability. Overcoming these limitations requires feedback-driven, closed-loop planning and control methods capable of adapting to the unpredictable and evolving conditions inherent in construction workflows.
@@
Learning-based methods have increasingly replaced traditional control-based methods for planning robot actions. Sensory–motor policy learning methods integrate sensor observations, planning, and control into a single end-to-end mapping, originally implemented using deep convolutional neural networks (CNNs) Levine et al. [2016]. These learned control policies map raw observations, such as images captured by a camera and force feedback, directly to motor commands or robot actions using various learning-based methods Finn et al. [2017], Levine et al. [2018], Kalashnikov et al. [2018], Brohan et al. [2023]. Among these methods, diffusion policy learning via behavior cloning Chi et al. [2024a] has demonstrated strong performance across various manipulation scenarios Chi et al. [2024b], Hou et al. [2025], Yang et al. [2025], Zhu et al. [2025]. The key advantages of diffusion policies lie in their ability to learn multimodal action distributions that capture multiple plausible ways to perform a task, while maintaining training stability and predicting coherent sequences of future actions Chi et al. [2024a].
@@
This paper addresses contact-rich robotic manipulation under fabrication uncertainty, defined here as the combined influence of tolerance accumulation across assembly steps and geometric variability arising from material imperfections, both of which introduce positional misalignment in tight-clearance timber joints and hinder precise and robust robotic assembly. We develop a sensory–motor diffusion policy learning framework via behavior cloning, trained on robot end effector pose and force/torque (F/T) data from teleoperated demonstrations, to plan and control the construction-scale assembly of mortise and tenon joints as a case study (Fig. 1).
The study pursues three primary objectives: 1) Demonstrate the applicability and performance of diffusion policies through a case study on the robotic assembly of mortise and tenon timber joints with sub-millimeter joint clearance; 2) Investigate policy robustness under fabrication uncertainties, simulated as randomized perturbations to the mortise position; and 3) Identify actionable guidelines for data collection strategies and policy parameter tuning that yield the most successful and robust performance, providing a practical reference for future research and deployment on construction-scale industrial robotic arms.
This paper presents one of the first systematic investigations in the AEC domain applying diffusion policy learning to robot planning and control. By enabling autonomous timber joinery assembly without reliance on human intervention, the proposed framework addresses physically demanding, craftsmanship-intensive tasks and offers a learning framework for high-precision contact-rich manipulation on large-scale industrial robots. The approach demonstrates its potential to generalize across a broad spectrum of contact-rich manipulation for assembly, including complex timber joints, pipe fittings, and light metal framing. Ultimately, this work advances robotic construction under uncertainty by bridging state-of-the-art sensory–motor policy learning with real-world AEC workflows and provides practical advances that can accelerate adoption across the field.
2 Related work
This section reviews prior work on robotic timber assembly, contact-rich manipulation, and sensory–motor policy learning, and learning-based approaches for construction applications. We first discuss the methods used for robotic timber assembly under uncertainty, including the contact-rich manipulation of timber joints. Next, we summarize recent advances in sensory–motor policy learning. Finally, we situate existing robot learning efforts in construction tasks, motivating the present study.
2.1 Robotic timber assembly
@@
In construction with discrete elements, such as brickwork and timber or steel framing, uncertainty remains a fundamental challenge to achieving precise and robust robotic assembly. In robotic timber assembly, which serves as the case study in this paper, these uncertainties arise from material imperfections (e.g., dimensional deviations, deformation, shrinkage/expansion due to moisture) and fabrication inaccuracies (e.g., cut length and angle errors, robot pose or grasp pose deviation). If unaccounted for, these perturbations cause error accumulation during assembly, leading to misalignment, part collision, and task failure when trajectories are executed directly from as-planned digital models  Adel et al. [2024]. Prior research in robotic timber assembly has explored various strategies to address these uncertainties, including computational modeling of tolerance propagation Gandia et al. [2022], adaptive, feedback-driven methods Helm et al. [2016], Adel et al. [2024], Ruan et al. [2023], Cote et al. [2024], and external tracking systems for end-effector pose correction in multi-robot assembly workflows Stadelmann et al. [2019], Adel [2020]. However, the successful and precise execution of timber joints often still depends on human intervention to correct deviations during nailing, gluing, and insertions Willmann et al. [2017], Adel et al. [2018], Adel [2020], Helmreich et al. [2022]. The challenges of uncertainty become even more pronounced in contact-sensitive assembly tasks, such as those in traditional timber joinery, where success depends not only on geometric precision but also on controlling excessive force interactions arising from friction and tight tolerances. In such scenarios, even minor deviations introduced during fabrication or caused by material imperfections can easily lead to misalignment, resulting in jamming, stick-slip behavior, and material damage Apolinarska et al. [2021], Kramberger et al. [2022].
2.2 Contact-rich manipulation of timber joints
@@
Handling force interactions in contact-rich robotic manipulation is often addressed through model-based impedance or admittance control laws Albu-Schäffer et al. [2007], Suárez-Ruiz et al. [2018], learning-based control policies Vecerik et al. [2019], Schoettler et al. [2020], or hybrid approaches combining both Johannsmeier et al. [2019], Hou et al. [2025]. In the context of assembly with timber joinery, a few studies have explored learning-based methods. Apolinarska et al. applied reinforcement learning for the motion control of an industrial robot to assemble single- and double-lap joints using robot pose and force/torque data from a wrist-mounted sensor Apolinarska et al. [2021]. For the single-lap joints, they also investigated policy robustness by introducing angular and linear offsets to the initial robot pose while grasping the top piece. Their results showed strong performance in zero-offset cases and revealed challenges of policy adaptation when angular and linear offsets, as well as angled or double insertions, were introduced. The findings underscored the importance of robust policy learning for handling geometric variations and positional deviations in contact-rich manipulation during assembly. Kramberger et al. also proposed a learning-by-demonstration method that integrated a compliance controller into Cartesian-space dynamic movement primitives for perpendicular single-lap joint insertion; however, they relied on significantly loose tolerances Kramberger et al. [2022].
@@
Other studies have explored semi-structured or open-loop methods for timber joinery assembly; Robeller et al. introduced a custom end-effector that generated vibration and combined it with a manually applied mallet-driven pulse force to facilitate the connection of wood panels using dovetail joints Robeller et al. [2017]. Leung et al. demonstrated the use of an industrial robot, combined with manually placed distributed robotic clamps, to assemble tight-fitting half-lap joints, highlighting the potential for precise, repeatable assembly Leung et al. [2021]. During their experiments, they observed challenges such as misalignments and collisions due to fabrication inaccuracies and deviations between digital and as-built models, which are common in construction-scale robotic applications. Rogeau employed visual feedback and fiducial markers for the assembly of wood panels with multiple mortise and tenon joints and reported successful full insertions in about 50% of their test cases across various joint configurations such as mortise tightness and tenon chamfer angles Rogeau [2023].
2.3 Sensory–motor policy learning
@@
Diffusion policies Chi et al. [2024a], Ze et al. [2024], Zhao et al. [2025] represent a recent class of behavior cloning methods built upon Denoising Diffusion Probabilistic Models (DDPMs) Ho et al. [2020], Sohl-Dickstein et al. [2015] and Denoising Diffusion Implicit Models (DDIMs) Song et al. [2022]. DDPMs are generative models originally developed for high-quality image synthesis, transforming random noise into samples from a target distribution through an iterative denoising process. DDIMs extend this framework by formulating a non-Markovian reverse process, which allows the number of denoising steps used during inference to be reduced without changing the training procedure, thereby enabling significantly faster sampling. In diffusion policy via robot action diffusion Chi et al. [2024a], the denoising process generates a sequence of robot trajectories at each inference step. This approach is known to capture multimodal action distributions, meaning it can represent multiple distinct yet equally valid action sequences for completing a task, as often occurs in human demonstrations. The approach also maintains temporal coherence, meaning the policy predicts an entire sequence of future actions jointly rather than one step at a time. This joint prediction ensures that consecutive actions remain consistent rather than switching erratically between different predicted single actions, which can lead to unstable or jittery motion. Moreover, the method avoids the need for action discretization (breaking continuous movements into a fixed set of bins), or negative sampling (generating incorrect actions during training to guide learning) Chi et al. [2024a]. This removes common limitations of explicit behavior cloning methods Florence et al. [2020], Shafiullah et al. [2022], which can make it difficult to achieve high-precision actions. Diffusion policies also offer improved training stability compared to implicit behavior cloning methods Florence et al. [2022], Jarrett et al. [2020]. In addition, they have demonstrated robustness to perturbations, visual occlusions, and idle actions in some manipulation tasks Chi et al. [2024a].
2.4 Robot learning in construction
Planning and control through contact remains a major challenge for both model- and learning-based approaches due to the inherent complexity of contact dynamics Yang et al. [2025]. While diffusion policies have achieved notable success in a range of dexterous manipulation tasks, their effectiveness in highly contact-rich scenarios is still under active investigation Hou et al. [2025], Kang et al. [2025], Wu et al. [2025], Zhou et al. [2026]. In particular, the deployment of these end-to-end sensory–motor policies on industrial manipulators for dexterous manipulation tasks has remained underexplored, likely due to compounded challenges arising from hardware constraints, high payloads, limited intrinsic compliance and force control, conservative safety-certified control interfaces, and restricted access to low-level control Kommey et al. [2025], Ruan et al. [2026], Delgado and Oyedele [2022].
A growing body of work has investigated robot learning for construction tasks, most commonly through imitation learning and reinforcement learning Sun et al. [2026], Huang and Zou [2026], Duan et al. [2024, 2025], Yu et al. [2024], Li and Zou [2023], Apolinarska et al. [2021]. Notably, most of these studies were conducted on lightweight collaborative or mobile manipulators operating under relatively compliant hardware conditions. Among these, Sun et al. investigated rebar slot insertion and rebar tying using a mobile manipulator, combining visual servoing for coarse positioning with imitation learning for precise execution Sun et al. [2026]. The rebar insertion task required precise alignment of the stirrup with three slots prior to insertion, with a reported total insertion clearance of approximately 1.5 mm. Similarly, Duan et al. studied a cable-in-duct installation task using a zero-shot Sim2Real visual–tactile reinforcement learning framework, evaluating robustness by varying pipe diameters (16–25 mm) for a 4 mm fish tape Duan et al. [2025].
Among learning-based studies in construction, Apolinarska et al. (mentioned in Section 2.2) is one of the few studies deploying reinforcement learning on a large-scale industrial manipulator, evaluating millimeter-scale (1 and 2 mm) geometric clearance for timber lap joint assembly Apolinarska et al. [2021]. Compared to these prior works, the present study investigates sensory–motor policy learning on industrial manipulators for timber joinery assembly under substantially tighter sub-millimeter clearances and positional perturbations that exceed these clearances, thereby probing the limits of policy robustness in a structurally constrained, contact-rich manipulation task.